OpenStudy (anonymous):

Express the complex number in trigonometric form. -4

OpenStudy (anonymous):

So, do you know how to find the the angle?

OpenStudy (anonymous):

Try to think of the unitcircle

OpenStudy (anonymous):

-4 is simply just a scaled-up -1 so that means that the angle is what?

OpenStudy (anonymous):

180 degrees?

OpenStudy (anonymous):

@frx

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

Or pi since you're looking for the polarform

OpenStudy (anonymous):

Thanks. These are the possible answers: 4(cos 270° + i sin 270°) 4(cos 90° + i sin 90°) 4(cos 0° + i sin 0°) 4(cos 180° + i sin 180°) So how do I find out which one of those is the answer if it is pi or 180 degrees? thanks again

OpenStudy (anonymous):

\[|z|=\sqrt{4^{2}}=4\] \[\arg(z)= \pi \] \[4(\cos \pi+isin \pi)\]

OpenStudy (anonymous):

So since your alternatives are in degrees and not radians you know it's 180 degrees.

OpenStudy (anonymous):

So it's the last one of your options

OpenStudy (anonymous):

Oh i see what you did. Thank you so much man. I really appreciate your help. Very thorough explanation.

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

If you have any questions just send a pm